(1) 4C(s) + 5H2(g)C4H10(g)...... ΔH° = -125.6 kJ

(2) C2H4(g)2C(s) + 2H2(g)......ΔH° = -52.3 kJ

what is the standard enthalpy change for the reaction:

(3) 2C2H4(g) + H2(g) --------> C4H10(g) ΔH° = ?

_______ kJ
1
Expert's answer
2019-04-04T07:08:42-0400

The standard enthalpy of formation of elementary substance is zero.


(1)         4C(s) + 5H2(g) = C4H10(g)...... ΔrH° = –125.6 kJ

ΔrH° = ΔfH°(C4H10) – 4 ΔfH°(C) – 5 ΔfH°(H2) = ΔfH°(C4H10) = –125.6 kJ.

ΔfH°(C4H10) = –125.6 kJ.


(2)        C2H4(g) = 2C(s) + 2H2(g)......ΔH° = –52.3 kJ

ΔrH° = 2ΔfH°(C) + 2 ΔfH°(H2) – ΔfH°(C2H4) = – ΔfH°(C2H4)  = –52.3 kJ

ΔfH°(C2H4)  = 52.3 kJ.


(3)        2C2H4(g) + H2(g) ----> C4H10(g)

ΔrH° = ΔfH°(C4H10) – 2ΔfH°(C2H4) – ΔfH°(H2) = ΔfH°(C4H10) – 2ΔfH°(C2H4) = –125.6 kJ – 2×52.3 kJ = –230.2 kJ.

ΔrH° = –230.2 kJ.